A medical lab is testing a new anticancer drug on cancer cells. The drug stock solution concentration is 1.5×10−91.5×10−9 M, and 1.00 mL of this solution will be delivered to a dish containing 2.0×1052.0×105 cancer cells in 5.00 mL of aqueous fluid. What is the ratio of drug molecules to the number of cancer cells in the dish?
Answer #1

Given drug stock solution concentration is 1.5×10−9 M; molarity = moles per liter (mol/L)

moles in 1 mL of this drug = 1 mL x 10-3 L x 1.5×10−9 M = 1.5 x 10-12 mol

1 mol of any substance will possess 6.02214129 × 1023 molecules/mol, so

molecules of drug present in 1.5 x 10-12 mol = 1.5 x 10-12 mol x 6.02214129 × 1023 molecules/mol = 9.033211935 x 1011 molecules.

ratio of drug molecules to the number of cancer cells in the dish = 9.033211935 x 1011 molecules / 2.0 ×105 cancer cells = 4.5166059675 ~ 4.52 x 106

ratio of drug molecules to the number of cancer cells in the dish = 4.52 x 106

Hope this helped you!

Thank You So Much! Please Rate this answer as you wish.("Thumbs Up")